show how to find the greatest common factor for:

42,90
34,68
56,150
28,45
60,350
40,68

Answer by checkley71(8403)

factors for 42=2*3*7
factors for 90=2*3*3*5 answer is 2*3=6
34=2*17
68=2*2*17 answer=2*17=34
56=2*2*14
150=2*3*5*5 answer=2
28=2*2*7
45=3*3*5 no common factors.
60=2*2*3*5
350=2*5*5*7 answer=2*5=10
40=2*2*2*5
68=2*2*17 answer=2*2=4